OpenStudy (anonymous):

how do I find the zeros for y= 12x^2+8x-15 and rewrite it in intercept form?

OpenStudy (anonymous):

factorize first

OpenStudy (anonymous):

y = 12x^2 + 8x - 15 dy/dx = m = 24x + 8 y = mx + c y = (24x + 8)x + c We need to find c, so: 12x^2 + 8x - 15 = (24x + 8)x + c 12x^2 + 8x - 15 = 24x^2 + 8x + c -12x^2 - 15 = c Therefore the form is: y = (24x + 8)x - (12x^2 + 15) Expand: y = 24x^2 + 8x - 12x^2 - 15 y = 12x^2 + 8x - 15 ... i dont know if i answered this correctly... was that your question??

OpenStudy (anonymous):

No, but thanks anyways. I have to have it look something like this. Ex: question: 2x^2+5x-3 answer: (2x-1)(x+3)

OpenStudy (anonymous):

oooohhh.... im sorry!! but u have the answer?

OpenStudy (anonymous):

Its fine. I have the answer to that question but I cannot figure out how to solve 12x^2+8x-15.

OpenStudy (anonymous):

u made it look sooo much easier!:)

OpenStudy (anonymous):

oh.. well i wish i could help, but i would probably just confuse you...:) sorry!!

OpenStudy (anonymous):

Its just factoring and the example was an easier problem. I have others similar to it. Ex: 18x^2+48+32 2(9x^2+24+16 2(3x+4)(3x+4) and the answer would be 2(3x+4)(3x+4)

OpenStudy (anonymous):

Thanks anyways though, Kelly.

OpenStudy (anonymous):

your welcome!